when the momentum of a body increases by 100% ,its kinetic energy is: pls. Explain.

Grade:12

3 Answers

Abhilash Kumar Prajapati
13 years ago

when momentum is increase by 100%,its kinetic energy is increase by 300%.

Mainak Chakraborty
13 years ago

As p direct. prop. v & KE direct. prop. v2 , as E incr. 100%, total v´ = 2v

So KE becomes 22 = 4 times,

And, % incr. in KE = 4E - E/E Χ 100 = 300 %

farhaan bukhsh
13 years ago
kinetic energy will become 3 times of original. p=mv 1/2mv*v=p*p/2m we can deduce it further